OpenStudy (anonymous):

log (x-4) + log (x+5)=1

OpenStudy (experimentx):

log((x-4)(x+5)) = log e (x-4)(x-5) = e

OpenStudy (anonymous):

But what does x equal?

jimthompson5910 (jim_thompson5910):

log (x-4) + log (x+5)=1 log[ (x-4)(x+5) ]=1 (x-4)(x+5) = 10^1 (x-4)(x+5) = 10 x^2+x-20 = 10 x^2+x-20 - 10 =0 x^2+x-30 =0 (x-5)(x+6) = 0 x-5=0 or x+6=0 x=5 or x=-6 Throw out the possible solution x = -6 since it's NOT in the domain. So the only answer is x = 5